OpenStudy (republican31):

(4n-6p)/p = pm/p for m

OpenStudy (republican31):

Please note I just corrected a mistake in it.

OpenStudy (anonymous):

What do u need to do?

OpenStudy (anonymous):

Solve for m?

OpenStudy (republican31):

Yes solve for M. Sorry for not responding.

OpenStudy (anonymous):

\[\frac{ (4n-6p) }{ p }=\frac{ pm }{ p }\]\[p(\frac{ (4n-6p )}{ p })=(\frac{ p m }{ p })p\]\[4n-6p=p m\]\[\frac{(4n-6p) }{ p }=m\] There's your answer. Do you understand that?

OpenStudy (anonymous):

From the 3rd step to the 4th step I divided the whole equation by p.

OpenStudy (republican31):

Yes, I think so. Just so you know the original problem was http://curriculum.kcdistancelearning.com/courses/ALG1x-FN-A08/a/exams/exam_E-Free_Response_Exam/img34.gif But I did what I could to simplify it, from when I asked it earlier and only got a partial response.

OpenStudy (anonymous):

You put pm/p

OpenStudy (anonymous):

But alright did you want me to solve this one for you or did you understand how to get it?

OpenStudy (republican31):

I understand it I think. It makes sense from the other concepts I've learned and been helped with in other questions. Is that still the answer for the original problem, in case I did something wrong originally?

OpenStudy (anonymous):

Which problem? The one from the link or the one I just did?

OpenStudy (anonymous):

What did you get for the problem you sent with the link? I'll check it for you.

OpenStudy (republican31):

The link. The link is the actual problem, from the beginning. But I tried to simplify it. I had asked this question with the original problem earlier, but people suddenly stopped responding without solving, so I took as far as they got and continued here. Maybe it's not correct.

OpenStudy (anonymous):

Oh wait I don't have to actually. You still get the same answer as the problem I showed you. :)

OpenStudy (republican31):

Yay! Thank you SOOOO much!

OpenStudy (anonymous):

Your welcome! :)